# MEMO: Dataset and Methods for Robust Multimodal Retinal Image Registration with Large or Small Vessel Density Differences

### v1

@@
Abstract
The measurement of retinal blood flow (RBF) in capillaries can provide a powerful biomarker for the early diagnosis and treatment of ocular diseases. However, no single modality can determine capillary flowrates with high precision. Combining erythrocyte-mediated angiography (EMA) with optical coherence tomography angiography (OCTA) has the potential to achieve this goal, as EMA can measure the absolute 2D RBF of retinal microvasculature and OCTA can provide the 3D structural images of capillaries. However, multimodal retinal image registration between these two modalities remains largely unexplored. To fill this gap, we establish MEMO, the first public multimodal EMA and OCTA retinal image dataset. A unique challenge in multimodal retinal image registration between these modalities is the relatively large difference in vessel density (VD). To address this challenge, we propose a segmentation-based deep-learning framework (VDD-Reg) and a new evaluation metric (MSD), which provide robust results despite differences in vessel density. VDD-Reg consists of a vessel segmentation module and a registration module. To train the vessel segmentation module, we further designed a two-stage semi-supervised learning framework (LVD-Seg) combining supervised and unsupervised losses. We demonstrate that VDD-Reg outperforms baseline methods quantitatively and qualitatively for cases of both small VD differences (using the CF-FA dataset) and large VD differences (using our MEMO dataset). Moreover, VDD-Reg requires as few as three annotated vessel segmentation masks to maintain its accuracy, demonstrating its feasibility.
{IEEEkeywords}
Erythrocyte mediated angiography, optical coherence tomography angiography, multi-modal retinal image registration.
@@ 1 Introduction
Current methods of measuring RBF are limited. For instance, laser Doppler imaging [6] is limited by high variability of measured flowrates. Dynamic OCTA [7, 8] and color Doppler imaging [9] can only measure relative flowrates, leading to poor intra-platform and cross-platform measurement repeatability. Adaptive optics scanning laser ophthalmoscopy (AO-SLO) [10, 11] and AO-OCT [12, 13] have a limited field of view. Erythrocyte mediated angiography (EMA) [14], on the other hand, is a novel technique which has the capability of determining absolute erythrocyte flowrates of arterioles and venules in vivo with high precision and a large field of view. EMA determines the flowrates by following the motion of individual fluorescently labelled erythrocyte ghosts in the retinal capillary circulation which can be visualized in vivo [15, 16, 17]. Despite the aforementioned advantages, a major limitation of EMA is that it cannot determine the axial location of capillaries. EMA images are captured in 2D using a scanning laser ophthalmoscope, but capillaries are located across different retinal layers [18].
@@
One potential solution to address the limitation of EMA is to combine it with other modalities providing 3D structural imaging of retinal capillaries. Optical coherence tomography angiography (OCTA) is an ideal candidate, as it can generate high-resolution 3D structural images of retinal capillaries [19, 20, 21]. Combining EMA and OCTA may enable absolute capillary RBF measurement for diagnosis and treatments of ocular diseases. A key requirement for this is accurate registration. Manual approaches to registration are time-consuming, necessitating the development of an automated approach to registration of EMA and OCTA image pairs.
@@
Multimodal retinal image registration has been extensively studied in recent years [22, 23, 24, 25, 26, 27, 28, 29]. However, current approaches have primarily utilized the public CF-FA dataset [30] (color fundus and fluorescein angiography) or private datasets with modalities other than EMA and OCTA, such as CF and fundus autofluorescence (FAF) [22, 23], and CF and infrared reflectance (IR) imaging [23, 27, 28]. The lack of new and publicly available multimodal retinal image datasets not only makes it difficult for researchers to fairly and thoroughly compare every method, but also prevents the identification of new challenges of registering image pairs from different modalities.
Figure 1: Sample images of (a) CF, (b) FA, (c) EMA and (d) OCTA with vessel density (VD). (a) and (b) are taken from the CF-FA dataset [30].
@@
To fill in this gap, we conducted experiments on nonhuman primates, to create a public dataset of EMA and OCTA pairs [31]. This dataset has the unique features of being well-controlled, one of the few datasets that includes OCtA images and the only dataset to include EMA sequences. We intend to expand the dataset to include human images. This dataset is described as the multimodal EMA and OCTA (MEMO) retinal image dataset. MEMO contains EMA and OCTA image pairs with manually labeled matched points for studying multimodal retinal image registration. Additionally, MEMO includes OCTA projection images [32] from three different layers (superficial vascular plexus (SVP), intermediate capillary plexus (ICP) and deep capillary plexus (DCP)) and EMA image sequences with moving erythrocytes.
@@
To overcome the challenges posed by large VD differences, we further propose VDD-Reg, a segmentation-based deep-learning framework for multimodal retinal image registration that can robustly register two imaging modalities despite vessel density differences. VDD-Reg consists of a vessel segmentation module and a registration module. For the vessel segmentation module, we also designed a novel two-stage semi-supervised learning framework, LVD-Seg, which requires only a few (e.g., three) labeled vessel segmentation masks from the modality with lower vessel density (EMA in our case). Specifically, LVD-Seg first uses a supervised loss (i.e., MSE) to stabilize the training of the vessel segmentation module, and then uses an unsupervised loss (i.e., style loss [33]) to further improve the segmentation and registration results.
Finally, we found that Soft Dice [27], a recently proposed image registration metric extended from Dice, could not accurately distinguish which method was better when large VD differences were present and when the results of each method did not exhibit significant differences. As a result, we propose a new registration metric called Masked Soft Dice (MSD). MSD considers only the pixels within the ground truth vessel segmentation masks of the modality with lower VD, which can more accurately evaluate each method in our MEMO dataset. With both MSD and standard registration metrics, we show that VDD-Reg can not only achieve the same level of accuracy as competing methods in datasets with similar vessel densities (CF-FA [30]), but also maintain high performance with large differences in VD, as in our MEMO dataset.
@@
    We establish MEMO, the first public multimodal EMA and OCTA retinal image dataset. MEMO provides registration ground truth, three layers of OCTA projection images and EMA image sequences containing moving erythrocytes. This has the potential for use in any resaerch involving OCtA registration with other modalities that use a scanning laser ophthalmoscope. MEMO is available at https://chiaoyiwang0424.github.io/MEMO/.
@@
    We address the unique challenge of retinal image registration between modalities with a large difference in vessel density (VD). To the best of our knowledge, MEMO is the first public multimodal retinal image dataset having this characteristic.
  3. 3.
@@
  4. 4.
    We introduce a registration metric, Masked Soft Dice (MSD), specifically designed for multimodal retinal image datasets with large VD differences.
@@
Dataset Multi-modal Modality Pair / Image Number Global Registration Ground truth
e-ophtha [34] - Color Fundus 144 pairs -
RODREP [35] - Color Fundus 1400 pairs -
VARIA [36] - Fundus Image 154 pairs -
FIRE [37] - Color Fundus 134 pairs
FLORI21 [38] - Ultra-widefield fluorescein angiography 15 pairs
CF-FA [30] 1. Color Fundus 2. Fluorescein Angiography 59 pairs -
@@ 2. Ultra-widefield fundus angiography
OCTA-500 [40] 1. OCTA 2. OCT 500 pairs -
MEMO (Our Dataset) 1. EMA 2. OCTA 30 pairs
2 Related Works
2.1 Retinal Image Datasets with Image Pairs
Existing public retinal image datasets with image pairs are listed in Table 1. Here, we focus on datasets with image pairs as they are more relevant to the main scope of this paper on multimodal retinal image registration. Private (in-house) datasets were excluded as they cannot be freely used to evaluate methods.
The datasets listed in Table 1 can be divided into monomodal and multimodal ones. Among the monomodal datasets, only FLORI21 [38] provides ultra-widefield fluorescein angiography images while the others [34, 35, 36, 37] provide fundus images. To be more specific, e-ophtha [34] contains 144 image pairs with large and small overlapping regions. RODREP [35] provides 1400 image pairs acquired from 140 eyes, but the overlap of each pair is limited. VARIA [36] provides 154 image pairs with only large overlapped areas due to its small FOV. The above three datasets do not provide a registration ground truth. In contrast, FIRE [37] provides registration ground truth, consisting of 134 fundus image pairs with both large and small areas of overlap. Ten corresponding points are provided for each image pair. FLORI21 [38] also provides registration ground truth and contains image pairs from 5 different subjects. Although FIRE and FLORI21 are useful for image registration research, they contain images from only one modality.
Table 1 demonstrates that public multimodal retinal image datasets with image pairs are relatively rare. CF-FA [30] contains 59 pairs of color fundus (CF) and fluorescein angiograms (FA). Twenty-nine pairs were collected from healthy eyes, while the rest were collected from eyes with diseases. Although CF-FA has been widely used for multimodal retinal image registration, the registration ground truth is not officially provided and the vessel density (VD) difference between the two modalities is relatively small. PRIME-FP20 [39] contains several ultra-wide field fundus photography (FP) and ultra-wide field fundus angiography (FA) retinal image pairs. However, it provides only the vessel segmentation ground truth of ultra-wide field fundus photography without providing the registration ground truth. OCTA-500 [40] provides 500 pairs of OCT and OCTA images with vessel segmentation ground truth. Despite its large size, the image pairs of OCT and OCTA were completely aligned, limiting its usage for multimodal retinal image registration.
Compared to the above datasets, MEMO has three major advantages. Firstly, MEMO is the first public multimodal retinal image dataset providing two modalities with relatively large VD differences. Secondly, the global registration ground truth is provided by manually labeling six corresponding point pairs per image pair. Finally, MEMO additionally provides raw EMA image sequences and OCTA projection images, which may be useful for multiple research fields such as automated erythrocyte tracking.
Figure 2: A sample EMA and OCTA pair from our MEMO dataset. Images inside the orange boxes were used for ground truth labeling. (A-1, A-2 and A-3: frame 0, 10 and 20 in the sample EMA image sequence. A-4: the stacked images of the EMA sequence. C-1, C-2 and C-3: the sample OCTA projection images representing DCP, ICP and SVP layer. C-4: the OCTA B-scan image. B and D: the six corresponding point pairs of the sample EMA and OCTA pair.)
@@ 2.2 Multi-Modal Retinal Image Registration
Multimodal retinal image registration methods can be categorized into conventional and deep learning-based methods. The conventional methods can be further divided into two types: direct and indirect methods. The direct conventional methods try to detect and match features directly on the raw images by manually designing more powerful feature descriptors or more robust matching algorithms. For example, Chen et al. [41] proposed a partial intensity invariant feature descriptor (PIIFD) and designed an image registration framework called Harris-PIIFD based on the proposed descriptor. Ghassabi et al. [42] combined UR-SIFT and PIIFD for image registration with large content or scale changes. Wang et al. [43] presented an image registration framework combining SURF, PIIFD and robust point matching. Lee et al. [44] introduced a low-dimensional step pattern analysis method to align retinal image pairs that were poorly aligned with baseline methods. Hossein-Nejad et al. [45] adopted adaptive Random Sample Consensus (A-RANSAC) for feature matching. On the other hand, the indirect conventional methods attempt to first transfer the images from different modalities into a similar ”style”, such as the vessel mask or the phase image, before detecting and matching features. For instance, Hernandez et al. [46] proposed line structures segmentation with a tensor-voting approach to improve registration. Hervella et al. [47] combined feature-based and intensity-based registration methods and employed a domain-adapted similarity metric to detect vessel bifurcations and crossovers. Motta et al. [48] proposed a registration framework based on optimal transport theory for vessel extraction on retinal fundus images. Li et al. [49] proposed a two-step registration method which converted raw images into phase images and adopted log-Gabor filters for global registration.
@@ 3.1 Overview
A sample EMA and OCTA image pair from the MEMO dataset is shown in Fig. 2. The dataset contains 30 pairs of EMA and OCTA images. For each image pair, 6 corresponding point pairs were manually annotated. The annotated points were chosen from the visually distinctive points in EMA and OCTA images, such as vessel bifurcation points and vessel bending points. All images were acquired following a protocol approved by Institutional Animal Care and Use Committee of the University of Maryland, Baltimore. Four eyes from two rhesus monkeys (macaca mulatta) were used to acquired paired EMA and OCTA images. Each pair was collected in the same session on the same date. Prior to the experimental session, the animal was sedated with ketamine and xylazine (5-10 and 0.2-0.4 mg/kg by intramuscular injection). The animal was intubated by trained veterinary technicians with an endotracheal tube and general anesthesia was maintained with 1.5% to 3% isoflurane with 100% oxygen. The animal was paralyzed with vecuronium (40-60 ug/kg, followed by 0.35-45 ug/kg/min), preventing eye movement during image acquisition. Body temperature was maintained at physiologic levels using a thermal blanket and blood pressure was monitored using a blood pressure cuff on the arm. The animal was laid in a prone position during the imaging session. A wire lid speculum was used to keep the eyelids open during imaging and tropicamide 1% was administered for pupillary dilation.
@@ 3.2 EMA
The procedure for EMA image acquisition is shown in Fig. 3. All EMA image sequences were captured by a Heidelberg Spectralis platform (Heidelberg Engineering, Heidelberg Germany). Approximately 17 mL of blood was drawn for processing with 5,6-carboxyfluorescein diacetate succinimidyl ester (CFSE) (Molecular Probes, USA) reconstituted in anhydrous dimethyl sulfoxide. Autologous erythrocytes were isolated from whole blood and loaded with 7.5 mM of CFSE using the osmotic shock method that has been detailed in prior publications [53]. Following cell preparation, up to 1.2 mL of CFSE-loaded cells were intravenously injected during image acquisition. After the cells were injected, ten-second angiograms centered on the macula were obtained with the Heidelberg Spectralis (Heidelberg Engineering GmbH, Germany) using a high speed 15-degree horizontal x 15-degree vertical field of view taken at 15 frames per second.
All image frames from the EMA image sequences were stored in TIF format. Six image sequences had the resolution of $512\times 512$ pixels, while the other 24 had the resolution of $384\times 384$ pixels. The pixel size of every EMA image sequence was provided. The stacked image of each EMA image sequence was used for registration ground truth labeling.

### v2

@@
MEMO: Dataset and Methods for Robust Multimodal Retinal Image Registration with Large or Small Vessel Density Differences
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
The measurement of retinal blood flow (RBF) in capillaries can provide a powerful biomarker for the early diagnosis and treatment of ocular diseases. However, no single modality can determine capillary flowrates with high precision. Combining erythrocyte-mediated angiography (EMA) with optical coherence tomography angiography (OCTA) has the potential to achieve this goal, as EMA can measure the absolute RBF of retinal microvasculature and OCTA can provide the structural images of capillaries. However, multimodal retinal image registration between these two modalities remains largely unexplored. To fill this gap, we establish MEMO, the first public multimodal EMA and OCTA retinal image dataset. A unique challenge in multimodal retinal image registration between these modalities is the relatively large difference in vessel density (VD). To address this challenge, we propose a segmentation-based deep-learning framework (VDD-Reg), which provides robust results despite differences in vessel density. VDD-Reg consists of a vessel segmentation module and a registration module. To train the vessel segmentation module, we further designed a two-stage semi-supervised learning framework (LVD-Seg) combining supervised and unsupervised losses. We demonstrate that VDD-Reg outperforms existing methods quantitatively and qualitatively for cases of both small VD differences (using the CF-FA dataset) and large VD differences (using our MEMO dataset). Moreover, VDD-Reg requires as few as three annotated vessel segmentation masks to maintain its accuracy, demonstrating its feasibility.
@@ 1 Introduction
Current methods of measuring RBF are limited. For instance, laser Doppler imaging [6] is limited by high variability of measured flowrates. Dynamic OCTA [7, 8] and color Doppler imaging [9] can only measure relative flowrates, leading to poor intra-platform and cross-platform measurement repeatability. Adaptive optics scanning laser ophthalmoscopy (AO-SLO) [10, 11] and AO-OCT [12, 13] have a limited field of view. Erythrocyte mediated angiography (EMA) [14], on the other hand, is a novel technique which has the capability of determining absolute erythrocyte flowrates of arterioles and venules in vivo with high precision and a large field of view. EMA determines the flowrates by following the motion of individual fluorescently labeled erythrocyte ghosts in the retinal capillary circulation which can be visualized in vivo [15, 16, 17]. Despite the aforementioned advantages, a major limitation of EMA is its inability to delineate the capillary structures through which the erythrocytes are flowing [18].
@@
One potential solution to address this limitation of EMA is to combine it with another modality that provides high-resolution structural imaging of retinal capillaries. Optical coherence tomography angiography (OCTA) is an ideal candidate, as it can generate high resolution imaging down to the capillary level in different layers of the retina [19, 20, 21]. Combining EMA and OCTA may enable absolute capillary RBF measurement for diagnosis and treatments of ocular diseases. A key requirement for this is accurate registration. Manual approaches to registration are time-consuming, necessitating the development of an automated approach to registration of EMA and OCTA image pairs.
@@
Multimodal retinal image registration has been extensively studied in recent years [22, 23, 24, 25, 26, 27, 28, 29]. However, current approaches have primarily utilized the public CF-FA dataset [30] (color fundus and fluorescein angiography) or private datasets with modalities other than EMA and OCTA, such as CF and fundus autofluorescence (FAF) [22, 23], and CF and infrared reflectance (IR) imaging [23, 27, 28]. The lack of new and publicly available multimodal retinal image datasets not only makes it difficult for researchers to fairly and thoroughly compare existing methods, but also prevents the identification of novel methods of multimodal imaging registration.
@@
To fill in these gaps in knowledge, we conducted experiments on non-human primates (NHPs) and created a public dataset of EMA and OCTA pairs. NHPs are used extensively in ophthalmic research and provide some of the best models for glaucoma as well as other ocular diseases [31]. The homology between NHPs and human eyes allows for ease in translation of experimental findings and applicability to human imaging and disease. Our dataset has the unique features of being well-controlled, one of the few datasets that includes OCTA images and the only dataset to include EMA sequences. This dataset is described as the multimodal EMA and OCTA (MEMO) retinal image dataset. MEMO contains EMA and OCTA image pairs with manually labeled matched points for studying multimodal retinal image registration. Additionally, MEMO includes OCTA projection images [32] from all three retinal vascular plexi (superficial vascular plexus (SVP), intermediate capillary plexus (ICP) and deep capillary plexus (DCP)) and EMA image sequences.
@@
Figure 1: Sample images of (a) CF, (b) FA, (c) EMA and (d) OCTA with vessel density (VD). (a) and (b) are taken from the CF-FA dataset [30]. In this example, the vessel density of OCTA (d) is five times grater than that of EMA (c) since most capillaries cannot be visualized in EMA images.
To overcome the challenges posed by large VD differences, we propose VDD-Reg, a segmentation-based deep-learning framework for multimodal retinal image registration that can robustly register two imaging modalities despite vessel density differences. VDD-Reg consists of a vessel segmentation module and a registration module. Here, instead of trying to extract every vessel as accurately as possible, the goal of the segmentation module is to extract vessels that are visible in both modalities so that the registration module can detect and match feature points more accurately. To achieve this goal, we designed a novel two-stage semi-supervised learning framework, LVD-Seg, which requires only a few (e.g., three) labeled vessel segmentation masks from the modality with lower vessel density (EMA in our case). Specifically, LVD-Seg first uses a supervised loss (i.e., MSE) to stabilize the training of the vessel segmentation module, and then uses an unsupervised loss (i.e., style loss [33]) with a unified style target image to guide the segmentation module to extract common vessels visible in both EMA and OCTA images, improving the registration accuracy.
@@
    We establish MEMO, the first public multimodal EMA and OCTA retinal image dataset. MEMO provides registration ground truth, all three retinal vascular plexi of OCTA projection images, and EMA image sequences containing moving erythrocytes. This also has the potential for use in any research involving OCTA registration with other modalities that use a scanning laser ophthalmoscope. MEMO is the first retinal image dataset containing EMA images and also the first multimodal retinal image registration dataset containing modalities with a large difference in vessel density (VD). MEMO is available at https://chiaoyiwang0424.github.io/MEMO/.
@@
2 Related Works
2.1 Retinal Image Datasets with Image Pairs
@@
Dataset Multi-modal Modality
Pair /
Image Number
Global Registration
Ground Truth
Initial Purpose
e-ophtha [34] - Color fundus 144 pairs - Diabetic retinopathy
VARIA [35] - Fundus image 154 pairs - User authentication
RODREP [36] - Color fundus 1400 pairs - Image registration
FIRE [37] - Color fundus 134 pairs Image registration
FLORI21 [38] - Ultra-widefield fluorescein angiography 15 pairs Image registration
OCTA-500 [39] 1. OCTA 2. OCT 500 pairs - Vessel segmentation
@@ 2. Ultra-widefield fundus angiography
15 pairs - Vessel segmentation
CF-FA [30]
1. Color fundus
2. Fluorescein angiography
@@
Diabetic retinopathy,
Image registration *
MEMO (Our Dataset) 1. EMA 2. OCTA 30 pairs Image registration
The registration ground truth is not officially provided.
There are relatively few public retinal image datasets specifically curated for image registration. In Table 1, we summarized the public retinal image datasets with image pairs as they could be potentially repurposed for image registration with proper ground truth annotations. The datasets listed in Table 1 can be divided into monomodal and multimodal. The monomodal datasets, such as e-ophtha [34], VARIA [35], RODREP [36], FIRE [37] and FLORI21 [38], contain images from only one modality, limiting their utility for multimodal retinal image registration research. On the other hand, existing multimodal retinal image datasets provide images from various modalities, such as OCT-OCTA pairs [39], ultra-widefield fundus photography-angiography pairs [40], and CF-FA pairs [30]. Compared to the above datasets, MEMO has three major advantages. Firstly, MEMO is the first retinal dataset with EMA images and also the first multimodal retinal image registration dataset providing two modalities with relatively large VD differences. Secondly, MEMO officially provides six corresponding point pairs per image pair as the global registration ground truth, which is crucial for fair comparisons of methods. Finally, MEMO provides raw EMA sequences and OCTA projection images, which may be useful for multiple research fields such as automated erythrocyte tracking.
@@ 2.2 Multi-Modal Retinal Image Registration
Multimodal retinal image registration methods can be categorized into conventional and deep learning-based methods. The conventional methods can be further divided into two types: direct and indirect methods. The direct conventional methods try to detect and match features directly on the raw images by manually designing more powerful feature descriptors or more robust matching algorithms. For example, Chen et al. [41] proposed a partial intensity invariant feature descriptor (PIIFD) and designed an image registration framework called Harris-PIIFD based on the proposed descriptor. Ghassabi et al. [42] combined UR-SIFT and PIIFD for image registration with large content or scale changes. Wang et al. [43] presented an image registration framework combining SURF, PIIFD and robust point matching. Lee et al. [44] introduced a low-dimensional step pattern analysis method to align retinal image pairs that were poorly aligned with baseline methods. Hossein-Nejad et al. [45] adopted adaptive Random Sample Consensus (A-RANSAC) for feature matching. On the other hand, the indirect conventional methods attempt to first transfer the images from different modalities into a similar "style", such as the vessel mask or the phase image, before detecting and matching features. For instance, Hernandez et al. [46] proposed line structures segmentation with a tensor-voting approach to improve registration. Hervella et al. [47] combined feature-based and intensity-based registration methods and employed a domain-adapted similarity metric to detect vessel bifurcations and crossovers. Motta et al. [48] proposed a registration framework based on optimal transport theory for vessel extraction on retinal fundus images. Li et al. [49] proposed a two-step registration method which converted raw images into phase images and adopted log-Gabor filters for global registration.
@@ 3.1 Overview
Figure 2: A typical sample EMA and OCTA pair from our MEMO dataset. Images inside the orange boxes were used for ground truth labeling. (A-1, A-2 and A-3: frame 0, 10 and 20 in the sample EMA image sequence. A-4: the stacked images of the EMA sequence. C-1, C-2 and C-3: the sample OCTA projection images representing DCP, ICP and SVP layer. C-4: the OCTA B-scan image. B and D: the six corresponding point pairs of the sample EMA and OCTA pair.)
A sample EMA and OCTA image pair from the MEMO dataset is shown in Fig. 2. The dataset contains 30 pairs of EMA and OCTA images. For each image pair, 6 corresponding point pairs were manually annotated. The annotated points were chosen from the visually distinctive points in EMA and OCTA images, such as vessel bifurcation points and vessel bending points. The procedure for EMA and OCTA image acquisition is shown in Fig. 3. All images were acquired following a protocol approved by Institutional Animal Care and Use Committee of the University of Maryland, Baltimore. Four eyes from two healthy non-human primates (rhesus monkeys, i.e., Macaca mulatta), aged 14 and 20, were used to acquired paired EMA and OCTA images. Each pair was collected in the same session on the same date. All image pairs were captured by a Heidelberg Spectralis platform (Heidelberg Engineering, Heidelberg Germany), which minimizes the nonlinear effect between two modalities or between different machines. Prior to the experimental session, the animal was sedated with ketamine and xylazine (5-10 and 0.2-0.4 mg/kg by intramuscular injection). The animal was intubated by trained veterinary technicians with an endotracheal tube and general anesthesia was maintained with 1.5% to 3% isoflurane with 100% oxygen. The animal was paralyzed with vecuronium (40-60 ug/kg, followed by 0.35-45 ug/kg/min), preventing eye movement during image acquisition. Body temperature was maintained at physiologic levels using a thermal blanket and blood pressure was monitored using a blood pressure cuff on the arm. The animal was laid in a prone position during the imaging session. A wire lid speculum was used to keep the eyelids open during imaging and tropicamide 1% was administered for pupillary dilation.
@@ 3.2 EMA
All EMA image sequences were captured by a Heidelberg Spectralis platform (Heidelberg Engineering, Heidelberg Germany). Each image sequence represents a time sequence capturing the trajectory of a single moving erythrocyte as it travels from the artery through the capillary to the vein. Approximately 17 mL of blood was drawn for processing with 5,6-carboxyfluorescein diacetate succinimidyl ester (CFSE) (Molecular Probes, USA) reconstituted in anhydrous dimethyl sulfoxide with a method similar to the human erythrocyte preparation, as previously described in our previous work [16]. Autologous erythrocytes were isolated from whole blood and loaded with 7.5 mM of CFSE using the osmotic shock method. Following cell preparation, up to 1.2 mL of CFSE-loaded cells were intravenously injected during image acquisition. After the cells were injected, ten-second angiograms centered on the macula were obtained with the Heidelberg Spectralis (Heidelberg Engineering GmbH, Germany) using a high speed 15-degree horizontal x 15-degree vertical field of view taken at 15 frames per second. More details of the procedure can be referred to our published protocol [16, 53, 54].
All image frames from the EMA image sequences were stored in TIF format. The number of image frames for each sequence is different, since the speed of each erythrocyte is different. In addition, six image sequences had the image size of $512\times 512$ pixels, while the other 24 had the image size of $384\times 384$ pixels. The reason for different sizes of image is because some images were taken in high-speed mode, while others were taken in high-resolution mode which has a higher pixel density. The pixel size of every EMA image sequence was provided. The stacked image of each EMA image sequence was used for registration ground truth labeling.